The area of a rectangle is x^2-x-12 square units. If the length is (x+3) units , then what is the perimeter of the rectangle
Answer by mananth(16946)

L * W
area = x^2-x-12
length = x+3
(x^2-x-12)/(x+3)
(x-4)(x+3)/(x+3)
= x+4 units is the width